# Verifix: Post-Training Correction to Improve Label Noise Robustness with Verified Samples

### v1

@@ Abstract
Label corruption, where training samples have incorrect labels, can significantly degrade the performance of machine learning models. This corruption often arises from non-expert labeling or adversarial attacks. Acquiring large, perfectly labeled datasets is costly, and retraining large models from scratch when a clean dataset becomes available is computationally expensive. To address this challenge, we propose Post-Training Correction, a new paradigm that adjusts model parameters after initial training to mitigate label noise, eliminating the need for retraining. We introduce Verifix, a novel Singular Value Decomposition (SVD) based algorithm that leverages a small, verified dataset to correct the model weights using a single update. Verifix uses SVD to estimate a Clean Activation Space and then projects the model’s weights onto this space to suppress activations corresponding to corrupted data. We demonstrate Verifix’s effectiveness on both synthetic and real-world label noise. Experiments on the CIFAR dataset with $25\%$ synthetic corruption show $7.36\%$ generalization improvements on average. Additionally, we observe generalization improvements of up to $2.63\%$ on naturally corrupted datasets like WebVision1.0 and Clothing1M. Find our code at https://github.com/sangamesh-kodge/Verifix.git.
Keywords:
Label Noise Robustness, Dataset Corruption, Singular Value Decomposition, Deep Neural Networks, Image Classification
1 Introduction
(a) Errors in Dataset Creation.
(b) Domain 1 - Dataset cleaning
(c) Domain 2 - Noise-Robust Training
(d) Proposed - Post-Training Correction
Figure 1: Overview of Label Noise Handling Components. (a) Illustrates the various sources of label noise - Intentional and Unintentional Errors, (b) Explores Data Cleaning approaches [30, 29, 10], (c) Presents Noise-Robust Training methods [9, 47], and (d) shows the proposed Post-Training Correction paradigm.
Deep learning has consistently showcased cutting-edge performance across a spectrum of learning tasks, spanning domains such as natural language processing and computer vision. These achievements largely stem from the substantial increase in available data, enabling more effective model training. However, the sheer size of these datasets poses challenges in ensuring accurate labeling. Recent studies such as [31], have revealed the prevalence of labeling errors in widely used benchmark datasets. These labeling errors can originate from various sources and fall into two broad categories: intentional and unintentional errors (Figure 0(a)). Intentional errors, often associated with data poisoning attacks[36], include label-flipping attacks where an adversary maliciously corrupts the training labels. These attacks aim to (1) compromise the model’s performance [1, 16], affecting the usability of the model for the desired application; (2) gain unauthorized access to circumvent the system of a safety-critical application through a backdoor attack [4]; or (3) render models more susceptible to membership inference attacks [5], unethically obtaining access to training data. Another source of label errors, called Unintentional errors, arises from the labeling process itself. For instance, crowd-sourcing and auto-labeling are inherently error-prone [35], and errors persist even when efforts are made to control them [22, 48]. While deep learning models exhibit robustness to a certain degree of label noise [33, 12], they are not immune to label noise, as demonstrated in [39].
To enhance label noise robustness, researchers have primarily concentrated on two research domains. Domain 1 (Figure 0(b)), focuses on dataset cleaning to ensure trustworthy model training by identifying and correcting mislabeled samples before learning. The naive approach of manual re-labeling by experts is infeasible for large datasets and also shown to be error-prone [35]. Hence, the majority of works focus on automated detection of corrupt samples [30, 29, 10]. However, many automated label corruption detection approaches leverage learning dynamics, incurring significant computational overhead. Domain 2 (Figure 0(c)) focuses on developing noise-robust algorithms to tackle label noise, this includes Sharpness-Aware Minimization (SAM)[9], MixUp[47], MentorMix[17], and dropout regularization [19]. Some of these approaches use compute-intensive training pipelines to counter the effect of label corruption in the training data. Despite these efforts, both domains have limitations: they can be computationally expensive, may not fully eliminate performance gaps compared to clean data training, and often necessitate complete retraining if label noise is discovered later. This presents a significant obstacle for large models trained on large datasets. To address the aforementioned challenges, we introduce a novel domain depicted in Figure 0(d)Post-Training Correction. This proposed domain focuses on adjusting model parameters after the training process, thereby eliminating the computational expense associated with full retraining.
In this work, we present Verifix, a novel Post-Training Correction algorithm that compensates for the effects of label noise using a small subset of verified clean training samples. Below, we outline the three essential components of Verifix (as depicted in Figure 2):
  • Trusted Data Estimation (Section 4.2) - Verifix leverages input curvature scores [10, 32] to automatically identify a subset of correctly labeled samples, $\mathcal{D}_{Trust}$, from a pre-trained model. Our emphasis on low-curvature samples yields up to a $1.06\%$ improvement in generalization scores compared to random sampling from clean training data.
  • Activation Alignment (Section 4.3) - Verifix performs Singular Value Decomposition (SVD) on the trusted dataset, $\mathcal{D}_{Trust}$, to decompose it into principal components. The left singular matrix ($U$) forms the basis directions of activation patterns associated with clean (trusted) data. We obtain the scaled basis vectors, where the basis directions $U$ are scaled according to the explained variance using the importance scaling $\Lambda$. This gives us the Activation Alignment Matrix.
  • Weight Update (Section 4.1) - Verifix leverages the Activation Alignment Space to update model weights, selectively reducing the influence of activations likely associated with corrupted data.
Figure 2: Overview of the proposed Verifix algorithm. Here, $\theta$ denotes the model weights which comprises of the layer weights represented by $W$.
We demonstrate the effectiveness of Verifix on both synthetically corrupted and real-world datasets. First, we present results for synthetically corrupted data by manually introducing noise into the CIFAR10/100 dataset [23]. When evaluating Verifix on ResNet18 [15] and VGG11 [37] network architectures, we observe significant improvement in generalization of $7.36\%$ under $25\%$ label corruption. Additionally, we test our algorithm on the Mini-WebVision [17], WebVision1.0[25], and Clothing1M [46] datasets, which contain naturally occurring noisy labels. On these real-world datasets, Verifix achieves up to $2.63\%$ generalization improvement on average despite the presence of natural label noise.
In summary, the contributions of this paper are:
  1. (i)
    We propose Post-Training Correction (Figure 0(d)), a label noise correction paradigm designed to improve model generalization. This paradigm focuses compute efficient post-training model parameter adjustments avoiding model retraining.
  2. (ii)
    We propose Verifix, a Singular Value Decomposition (SVD) based technique, which establishes a new baseline to improve the generalization performance of the model trained with noisy labels Post-Training. Verifix automates the selection of trusted samples using curvature scores and updates model weights in a single step, offering significant computational savings over standard retraining approaches.
  3. (iii)
    We empirically validate the performance of our method for both synthetically and naturally corrupted datasets, and show a generalization improvement up to $10.19\%$ (for $25\%$ corruption) and $2.63\%$, respectively.
2 Related Work
We survey existing works in two key domains of label noise research:
2.0.1 Data Cleaning (Domain 1) :
This direction of research focuses on detecting and correcting or removing mislabeled samples within the training dataset, ensuring the model only learns from accurately labeled data. The naive human (expert) in the loop relabeling is tedious and often error-prone [35]. In response, the research in this domain primarily focuses on the automatic detection of mislabeled samples. Several works [8, 2, 41, 14, 38, 50, 20] train an ensemble of models to filter out the incorrectly labeled samples. A major limitation of these ensemble-based approaches is their computational expense, especially for large datasets where training multiple models can be infeasible. To address this challenge, several studies have focused on leveraging training dynamics to identify label noise, including methods such as using cross-validation on noisy splits [3], learning time [26, 45], forgetting time [29], train-set confidence [30] or thorough input loss curvature [10, 32]. While these approaches reduce computational overheads compared to ensemble-based methods, a remaining challenge lies in distinguishing between truly difficult examples and those that are mislabeled [10]. Additionally, approaches in this domain still necessitate retraining models on the filtered dataset, failing to fully leverage the models learned on the initial noisy data. In contrast, our proposed Post-Training Correction paradigm seeks computationally efficient ways to directly adjust the corrupted model. This approach is crucial for large datasets and models where retraining might be computationally intractable.
2.0.2 Noise-Robust Training (Domain 2) :
This domain of algorithms tackles the challenge of label noise through the development of training techniques and loss functions that are inherently less sensitive to mislabeled data. Regularization techniques, such as dropout [19] (which introduces randomness during training) and label smoothing [28, 44] (which softens target labels), help prevent overfitting and improve label noise robustness. Robust loss functions like symmetric cross-entropy [43] and Mean Absolute Error (MAE) reduce the impact of noisy labels compared to standard cross-entropy loss. Optimization methods like Sharpness-Aware Minimization (SAM) [9] promote flatter loss landscapes, which are associated with better generalization under label noise. Additionally, curriculum and self-paced learning [18, 49] gradually increase the difficulty of training examples, helping the model learn robust features before encountering noise. Data augmentation and mixup techniques (e.g., MixUp [47], DivideMix [24], MentorMix[17]) smooth training distributions, encouraging less abrupt decision boundaries and further enhancing robustness. While these robust algorithms offer improvement over standard training, they still have limitations. Some can be computationally expensive to implement or may reduce performance on clean datasets. Additionally, they often underperform compared to models trained with perfectly labeled data. Our proposed algorithm is orthogonal to these techniques and can be used in combination to further close this generalization gap.
3 Background
3.1 Singular Value Decomposition
@@
A rectangular matrix $A$ in $\mathbb{R}^{d\times n}$ can be decomposed with Singular Value Decomposition (SVD) [7] as
@@
3.2 Problem Formulation
Let the training dataset be defined as $\mathcal{D}_{Tr}=\{(x_{i},y_{i})\}_{i=1}^{N_{Tr}}$, where $N_{Tr}$ represents the total number of training samples. Here, $x_{i}$ denotes the input features with the corresponding label $y_{i}$. The validation dataset is represented as $\mathcal{D}_{Val}=\{(x_{i},y_{i})\}_{i=1}^{N_{Val}}$, and the test dataset is denoted as $\mathcal{D}_{Te}=\{(x_{i},y_{i})\}_{i=1}^{N_{Te}}$, where $N_{Val}$ and $N_{Te}$ signify the number of validation samples and test samples, respectively. Consider a deep neural network that approximates the mapping from input data $x_{i}$ to the corresponding label $y_{i}$ through the function $y=f(\theta,x_{i})$, where $\theta$ denotes the network’s weights or parameters. These parameters are trained through gradient descent-based optimization on the samples from the training dataset $\mathcal{D}_{Tr}$, and the validation dataset $\mathcal{D}_{Val}$ is used for tuning the hyperparameters. For a network with strong generalization performance, we expect that the relationship $y_{i}=f(\theta,x_{i})$ holds true for a substantial number of samples $(x_{i},y_{i})\in\mathcal{D}_{Te}$. Since network parameters are trained to minimize the empirical risk or loss $\mathcal{L}(y_{i},f(\theta,x_{i}))$ on the samples from $\mathcal{D}_{Tr}$, the presence of label corruption or noise in a portion of the training dataset can hamper network’s generalization ability thus decreasing the performance.
Given : $\begin{split}\text{Given : }&\theta_{*}=\text{{train}}(\mathcal{D}_{Tr})\\ \text{Obtain : }&\hat{\theta}_{*}=\text{{post\_training\_correction}}(\theta_{*},\mathcal{D}_{Tr})\\ \text{ s.t. }&y_{i}\approx f(\hat{\theta}_{*},x_{i})\approx f(\theta_{*}^{cln},x_{i})\\ \text{ where }&(x_{i},y_{i})\in\mathcal{D}_{Val}\text{ and }\theta_{*}^{cln}=\text{{train}}(\mathcal{D}_{Tr}^{cln})\end{split}$ (3)
The objective of a label noise correction algorithm is to enhance the generalization performance of a network trained on data, $\mathcal{D}_{Tr}$ containing noisy/corrupted labels. This essentially involves removing the influence of corrupted samples from the network parameters $\theta_{*}$ trained on corrupted data. Such an algorithm is expected to produce parameters $\hat{\theta}_{*}$ that closely matches the functionality of the model with parameters $\theta_{*}^{cln}$, which were trained on $\mathcal{D}_{Tr}^{cln}\subset\mathcal{D}_{Tr}$, a subset comprising correctly labeled samples. This objective be formally expressed as Equation 3.
4 Verifix
A network trained on a dataset $\mathcal{D}_{Tr}$ with corrupted labels is prone to overfit on the incorrect samples within the training data. Such a network is likely to produce spurious intermediate activations to incorporate the incorrect sample-label mapping. Our label correction approach (Figure 2), outlined in Algorithm 1, aims to align these intermediate activations with representative or trusted activations. In the next Subsection, we explain our weight update mechanism which enables us to modify the model weights to remove the influence of corrupt data. 111 We provide our code at https://github.com/sangamesh-kodge/Verifix.git.
4.1 Weight Update
@@
procedure post_training_correction( $\theta_{*}$, $\mathcal{D}_{Tr}$, $\alpha$)
1. $\mathcal{D}_{Trust}$ = get_trusted($\theta_{*}$, $\mathcal{D}_{Tr}$) // Section 4.2
2. $R$ = representation($\theta_{*}$, $\mathcal{D}_{Trust}$) // Eqn. 6, 7
@@
5.     $\Lambda^{l}$ = scale_importance($\Sigma^{l}$, $\alpha$ ) // Eqn. 8
6.     $W_{p}^{l}$ = $U^{l}\Lambda^{l}(U^{l})^{T}$ // Eqn. 9
7.     $\widehat{W}^{l}_{*}=$ update_parameter($W_{p}^{l}$, $W^{l}_{*}$ ) // Eqn. 4
@@
Algorithm 1 Verifix algorithm.
Obtaining a high-quality alignment matrix requires access to representative activations for ‘trusted’ samples or the samples in $\mathcal{D}_{Tr}$ with correct labels. We, therefore, divide the step of obtaining $W_{p}$ into two parts: (1) estimating a few correctly labeled or ‘trusted’ samples from $\mathcal{D}_{Tr}$, as presented in Section 4.2, and (2) obtaining the Activation Alignment Matrix using these samples outlined in Section 4.3.
4.2 Trusted Data Estimation
We define the trusted dataset, $\mathcal{D}_{Trust}=\{(x_{i},y_{i})\}_{i=1}^{N_{Trust}}$, as a subset of $\mathcal{D}_{Tr}$ containing a few ($N_{Trust}$) correctly labeled samples. This can be obtained by employing human expert labelers. Instead, in this paper we propose an automatic approach using the curvature scores [11, 10] of the training samples to identify samples that are correctly labeled. Note we use the curvature at the last epoch rather than the average curvature over all the training epochs [32]. This score is given by Equation 5
$Curv(x_{i},y_{i})=\frac{1}{n}\sum_{j=0}^{n}\left\lVert\frac{\partial(L(x_{i}+hv_{j},y_{i})-L(x_{i},y_{i}))}{\partial x_{i}}\right\rVert_{2}^{2}$ (5)
Where $n$ is the number of Rademacher variables $v_{j}$ to average over. Using low curvature input samples is a good way to obtain trusted data. It was shown in [10, 11] that curvature performs well when compared to other mislabeling detection methods since it captures network memorization [32]. We compute curvature scores for all the samples in $\mathcal{D}_{Tr}$ and select the $N_{Trust}$ samples with the lowest curvature scores to obtain the trusted datasets $\mathcal{D}_{Trust}$. These samples are used to obtain the Activation Alignment Matrix in the next Subsection.
4.3 Activation Alignment
@@
4.3.1 Representation Sampling :
For the $l^{th}$ linear or convolutional layer of the network, we collect the input activation $a_{in}$ and store these activation in a representation matrix $R^{l}$. This matrix captures representative information from the trusted samples in the input activations. It is utilized to estimate the trusted activation basis. Next, we elaborate on the details of this representation matrix [34] for each type of layer.
  1. (a)
@@ 4.3.1 Representation Sampling :
  2. (b)
    Convolutional Layer - For the convolutional layer, we express the convolution operation as matrix multiplication to apply the weight update rule proposed in Equation 4. This is achieved through the unfold [27] operation where the convolution operation is represented as matrix multiplication. Let the kernel size of the convolutional layer be $C_{out}\times C_{in}\times k\times k$, where $C_{in}$ represents the number of input channels, $C_{out}$ denotes the number of output channels and $k$ is the kernel size. The unfold operation would flatten all the patches in the input activations on which this kernel operates in a sliding window fashion, which gives us a matrix of size $n_{p}\times C_{in}kk$. Here, $n_{p}$ represents the number of patches in the activation $a_{in}^{i}$ of the $i^{th}$ sample in $\mathcal{D}_{Trust}$. This process allows us to represent convolution as matrix multiplication between the reshaped weights of size $C_{out}\times{C_{in}kk}$ and the unfolded activation of $a^{i}_{in}$ of size $n_{p}\times{C_{in}kk}$. Figure 9 in Appendix details the conversion of the convolution operation into the matrix multiplication operation. The unfolded activations of all the samples are concatenated and stored in the representation matrix as represented by Equation 7. We use the reshaped convolutions parameters to update the parameters as given in Equation 4.
@@ 4.3.2 SVD on Representations :
4.3.3 Importance Scaling :
We formulate a diagonal importance matrix $\Lambda$ with the $j^{th}$ diagonal component $\lambda_{j}$ given by Equation 8 [21]. Here, $\widetilde{\sigma}_{j}$ represents the normalized $j^{th}$ singular value, as defined in Equation 2. The effect of such scaling on singular values is shown in Figure 3. The parameter $\alpha\in(0,\infty)$, known as the scaling coefficient, serves as a hyperparameter controlling the scaling of the basis vectors. When $\alpha$ is set to 1, the basis vectors are scaled by the amount of variance they explain. Figure 2(b) shows the effect of changing $\alpha$ on the scaling of the basis vectors. As $\alpha$ increases, the importance score for each basis vector increases and approaches 1 as $\alpha\to\infty$. Conversely, a decrease in $\alpha$ diminishes the importance of the basis vectors, approaching 0 as $\alpha\to 0$. The scale_parameter procedure in line 5 of Algorithm 1 corresponds to this importance-based scaling.

### v2

# SAP: Corrective Machine Unlearning with Scaled Activation Projection for Label Noise Robustness

@@ Abstract
Label corruption, where training samples are mislabeled due to non-expert annotation or adversarial attacks, significantly degrades model performance. Acquiring large, perfectly labeled datasets is costly, and retraining models from scratch is computationally expensive. To address this, we introduce Scaled Activation Projection (SAP), a novel SVD (Singular Value Decomposition)-based corrective machine unlearning algorithm. SAP mitigates label noise by identifying a small subset of trusted samples using cross-entropy loss and projecting model weights onto a clean activation space estimated using SVD on these trusted samples. This process suppresses the noise introduced in activations due to the mislabeled samples. In our experiments, we demonstrate SAP’s effectiveness on synthetic noise with different settings and real-world label noise. SAP applied to the CIFAR dataset with $25\%$ synthetic corruption show upto $6\%$ generalization improvements. Additionally, SAP can improve the generalization over noise robust training approaches on CIFAR dataset by $\sim 3.2\%$ on average. Further, we observe generalization improvements of $2.31\%$ for a Vision Transformer model trained on naturally corrupted Clothing1M.
Codehttps://github.com/sangamesh-kodge/SAP.git
Introduction
Deep learning models have revolutionized various fields like natural language processing and computer vision, largely due to the availability of massive datasets. However, the very scale of these datasets presents a challenge: guaranteeing accurate labeling as shown in Figure 1. Recent studies (Northcutt, Athalye, and Mueller 2021), have exposed the significant presence of labeling errors in widely used benchmarks. These errors can be unintentional, arising from human error (Sambasivan et al. 2021) or in recent times automated labeling using Large Language models (Wang et al. 2024; Pangakis and Wolken 2024). Alternatively, they can be deliberate, introduced through attacks like data poisoning (Schwarzschild et al. 2021; Biggio, Nelson, and Laskov 2012; Jagielski et al. 2018; Chen et al. 2017, 2022). Regardless of origin, these errors degrade model performance. While deep learning models exhibit robustness to a certain degree of label noise, they are not immune to label noise, as demonstrated in (Steinhardt, Koh, and Liang 2017).
Figure 1: Sources of label noise - Intentional/Unintentional Errors.
Corrective Machine Unlearning (Goel et al. 2024) is emerging as a promising solution to the aforementioned challenge, with the objective of reducing the impact of mislabeled data by manipulating the trained model. A vast majority of unlearning algorithms (Triantafillou et al. 2024) require knowledge of which samples are mislabeled to partition the data into the retain set and the forget set. It is challenging to distinguish between mislabeled samples and hard to learn samples (Garg, Ravikumar, and Roy 2024). This presents a major challenge to deploy machine unlearning algorithms for tackling label noise.
This work introduces SAP, a novel algorithm based on Singular Value Decomposition (SVD) to improve model generalization in the presence of label noise. Notably, our approach achieves this improvement with a single update to the model weights. SAP leverages a small set of samples identified as having low cross-entropy loss. These samples are hypothesized to be correctly labeled and form a “clean” subset. We utilize this clean subset and SVD to compute an update for the model weights. This update involves estimating a clean activation space and projecting the model weights onto this space. This projection aims to suppress activations corresponding to potentially corrupted activations, improving model performance on unseen data. As SAP relies on the estimated clean subset (or retain set), we avoid explicitly detecting mislabeled data, which can be more difficult to obtain.
In order to visualize the effectiveness of SAP, we trained a network for binary classification task on clean training data (clean model) and corrupt data (corrupted model), where $10\%$ labels were flipped from 2D spiral dataset (Verma et al. 2019). We present a detailed experimental details for this toy illustration in the Supplementary. Figure 2(a) and 2(b) show the decision boundary of the model for the former and later cases respectively. We observe that, with label corruption, the test accuracy of the model drops by $4.8\%$ from the model trained without any label corruption. The model trained on noisy data in Figure 2(b) has a more complex and rough decision boundary which indicates memorization of the corrupt data points leading to a loss in generalization. When SAP is applied to the model trained on corrupt data as seen in Figure 2(c), we observe improved decision boundary smoothness and recover the generalization performance as hypothesized.
(a) Clean model
(b) Corrupt model
(c) SAP applied to corrupt model
Figure 2: Decision boundaries for a network trained the 2D spiral data (a) with full Clean data, (b) with $10\%$ corrupt data and (c) SAP applied to network trained on corrupt data.
We summarize the contributions of our work below:
  • We propose SAP, a novel corrective unlearning algorithm based on Singular Value Decomposition (SVD). SAP automates the selection of clean samples using the model loss and updates model weights in a single step, leading to computational efficiency. Notably, we alleviate the challenge of explicitly detecting mislabeled data which is required by most of the unlearning algorithms.
  • We empirically validate SAP’s performance on synthetic and real-world label noise scenarios across various model architectures and datasets, demonstrating generalization improvements of up to $6\%$ and $2.31\%$, respectively.
Background
Singular Value Decomposition
@@ Introduction
A rectangular matrix $A$ in $\mathbb{R}^{d\times n}$ can be decomposed with Singular Value Decomposition (SVD) (Deisenroth, Faisal, and Ong 2020) as
@@
Related Work
Corrective Machine Unlearning (Goel et al. 2024) has emerged as a promising approach to tackle the challenge of mitigating the impact of corrupted data, such as mislabeled samples, on trained models. Unlike traditional machine unlearning (Triantafillou et al. 2024), which often focuses on privacy concerns when removing data, corrective unlearning prioritizes improving model generalization, bypassing the need for adhering to specific privacy requirements (Hayes et al. 2024). In the context of label noise, corrective unlearning aims to improve the model’s generalization on unseen data, when the training dataset contains mislabeled samples. Recent work by (Goel et al. 2024) benchmarks several state-of-the-art (SoTA) unlearning algorithms, such as SSD (Foster, Schoepf, and Brintrup 2024), CF-k (Goel et al. 2022) and SCRUB (Kurmanji, Triantafillou, and Triantafillou 2023), within the corrective unlearning framework. Notably, ASSD (Schoepf, Foster, and Brintrup 2024) proposes an extension of SSD which automatically chooses the suitable unlearning hyperparameters to handle the label errors on supply chain delay prediction problem.
Additionally, model generalization can be further improved by using these algorithms in conjunction with traditional approaches such as data filtering (Northcutt, Jiang, and Chuang 2021; Jia et al. 2022; Maini et al. 2022), sample selection (Cheng et al. 2021), label correction (Zheng, Awadallah, and Dumais 2021), regularization (Wei et al. 2021, 2023), robust loss functions (Wang et al. 2019), optimizations (Foret et al. 2021), curriculum learning (Jiang et al. 2018; Zhang and Sabuncu 2018), data augmentation (Zhang et al. 2017; Jiang et al. 2020), and noise transition matrix estimation (Zhu, Wang, and Liu 2022; Cheng et al. 2022). We provide a brief Literature Survey on Label Noise Learning in the supplementary.
Scaled Activation Projection Algorithm
Figure 3: Overview of the proposed SAP algorithm. Here, $\theta_{*}$ denotes the model weights which comprises of the layer weights represented by $W$.
A network trained on the training dataset, $\mathcal{D}_{Tr}$, with corrupted labels is prone to overfit on the incorrect samples within the training data, as seen in Figure 2(b). Such a network is likely to produce spurious intermediate activations to incorporate the incorrect sample-label mapping. SAP (Figure 3), outlined in Algorithm 1, aims to align these intermediate activations with representative or clean activations. In the next Subsection, we explain our weight update mechanism which enables us to modify the model weights to remove the influence of corrupt data. We provide our codebase at https://github.com/sangamesh-kodge/SAP.git.
Weight Update
@@
procedure SAP( $\theta_{*}$, $\mathcal{D}_{Tr}$, $\alpha$)
1. $\mathcal{D}_{Trust}$ = get_trusted($\theta_{*}$, $\mathcal{D}_{Tr}$) // Eqn 4
2. $R$ = representation($\theta_{*}$, $\mathcal{D}_{Trust}$) // Eqn. 5, 6
@@ 1. $\mathcal{D}_{Trust}$ = get_trusted($\theta_{*}$, $\mathcal{D}_{Tr}$) // Eqn 4
5.     $\Lambda^{l}$ = scale_importance($\Sigma^{l}$, $\alpha$ ) // Eqn. 7
6.     $W_{p}^{l}$ = $U^{l}\Lambda^{l}(U^{l})^{T}$ // Eqn. 8
7.     $\widehat{W}^{l}_{*}=$ update_parameter($W_{p}^{l}$, $W^{l}_{*}$ ) // Eqn. 3
@@
Algorithm 1 SAP algorithm.
Obtaining a high-quality alignment matrix requires access to representative activations for ‘trusted’ samples or the samples in $\mathcal{D}_{Tr}$ with correct labels. We, therefore, divide the step of obtaining $W_{p}$ into two parts: (1) approximately estimating a few correctly labeled or ‘trusted’ samples from $\mathcal{D}_{Tr}$, and (2) obtaining the Activation Alignment Matrix using these samples.
Trusted Data Estimation
We define the trusted dataset, $\mathcal{D}_{Trust}=\{(x_{i},y_{i})\}_{i=1}^{N_{Trust}}$, as a subset of $\mathcal{D}_{Tr}$ containing a few ($N_{Trust}$) correctly labeled samples. This work proposes a method to automate the process of identifying these trusted samples. We leverage the cross-entropy loss, denoted by $\mathcal{L}$, of the trained model with parameters $\theta_{*}$. The rationale is that samples with lower cross-entropy loss are more likely to be correctly labeled. We select $N_{Trust}$ samples with the lowest cross-entropy loss values to from the trusted dataset $\mathcal{D}_{Trust}$. This process can be mathematically expressed by Equation 4 corresponds to get_trusted of Algorithm 1. Note, these trusted samples can also be obtained by employing human experts, however, this might be tedious or impractical for larger dataset.
$\begin{split}\mathcal{D}_{Trust}&=\operatorname*{argmin}_{\mathcal{S}}\sum_{(x_{i},y_{i})\in S}(\mathcal{L}(\theta_{*},x_{i},y_{i}));\\ &\text{such that }\mathcal{S}:=\{\mathcal{S}_{i}\subseteq\mathcal{D}_{Tr}\hskip 5.69054pt|\hskip 5.69054pt|\mathcal{S}_{i}|=N_{Trust}\}\end{split}$ (4)
Activation Alignment
@@
Representation Sampling:
For the $l^{th}$ linear or convolutional layer of the network, we collect the input activation $a_{in}$ and store these activation in a representation matrix $R^{l}$. This matrix captures representative information from the trusted samples in the input activations. It is utilized to estimate the trusted activation basis. Next, we elaborate on the details of this representation matrix (Saha, Garg, and Roy 2021) for each type of layer.
  1. 1.
@@
  2. 2.
    Convolutional Layer - For the convolutional layer, we express the convolution operation as matrix multiplication to apply the weight update rule proposed in Equation 3. This is achieved through the unfold (Chetlur et al. 2014) operation where the convolution operation is represented as matrix multiplication. Let the kernel size of the convolutional layer be $C_{out}\times C_{in}\times k\times k$, where $C_{in}$ represents the number of input channels, $C_{out}$ denotes the number of output channels and $k$ is the kernel size. The unfold operation would flatten all the patches in the input activations on which this kernel operates in a sliding window fashion, which gives us a matrix of size $n_{p}\times C_{in}kk$. Here, $n_{p}$ represents the number of patches in the activation $a_{in}^{i}$ of the $i^{th}$ sample in $\mathcal{D}_{Trust}$. This process allows us to represent convolution as matrix multiplication between the reshaped weights of size $C_{out}\times{C_{in}kk}$ and the unfolded activation of $a^{i}_{in}$ of size $n_{p}\times{C_{in}kk}$. Figure 8 in supplementary details the conversion of the convolution operation into the matrix multiplication operation. The unfolded activations of all the samples are concatenated and stored in the representation matrix as represented by Equation 6. We use the reshaped convolutions parameters to update the parameters as given in Equation 3.
@@
Importance Scaling:
We formulate a diagonal importance matrix $\Lambda$ with the $j^{th}$ diagonal component $\lambda_{j}$ given by Equation 7 (Kodge, Saha, and Roy 2024). Here, $\widetilde{\sigma}_{j}$ represents the normalized $j^{th}$ singular value, as defined in Equation 2. The effect of such scaling on singular values is shown in Figure 4. The parameter $\alpha\in(0,\infty)$, known as the scaling coefficient, serves as a hyperparameter controlling the scaling of the basis vectors. When $\alpha$ is set to 1, the basis vectors are scaled by the amount of variance they explain. Figure 4(b) shows the effect of changing $\alpha$ on the scaling of the basis vectors. As $\alpha$ increases, the importance score for each basis vector increases and approaches 1 as $\alpha\to\infty$. Conversely, a decrease in $\alpha$ diminishes the importance of the basis vectors, approaching 0 as $\alpha\to 0$. The scale_parameter procedure in line 5 of Algorithm 1 corresponds to this importance-based scaling.